A person has been diagnosed with third degree burns over the anterior surfaces of their arms and trunk. What specific structural damage has occurred to their skin and what risks to life may result from this injury? Use the Rule of Nines to estimate the percentage of burned skin. Please type it out
Answer #1

Answer: Third-degree burns are full thickness burs. It causes destruction of epidermis, dermis as well as the innermost subcutaneous tissues.

Risks to life:

Loss of Keratin ad Langerhans cells increases the risk of microbial invasion.

Loss of keratinized structure ad lamellar granules causes excessive loss of water, plasma, plasma proteins which leads to shock. This leads to dehydration, decreased blood volume and circulation, decreased urine output and diminished immune responses.

Percentage of burned skin using Rule of nines.

Here the body surface area o an adult who is burned is calculated using multiples of 9.

Anterior of arm= 4.5%

For both arms=2 x 4.5=9%

Trunk=9+9=18%

TOTAL=9+18 = 27%